# Guided Discrete Diffusion for Electronic Health Record Generation

### v1

@@ 5.1 Experiment Setup
Datasets. Public Datasets MIMIC-III (Johnson et al., 2016) includes deidentified patient EHRs from hospital stays. For each patient’s EHR, we extract the diagnosis and procedure ICD-9 codes and truncate the codes to the first three digits. This dataset includes a patient population of size 46,520. Private Datasets We consider two private datasets of patient EHRs from a large healthcare institution. For each, we extract the diagnosis and procedure ICD-10 codes and truncate the codes to the first three digits. The first dataset, denoted by $\mathcal{D}_{1}$, includes a patient population of size 1,670,347 and has sparse binary features; the second dataset, denoted by $\mathcal{D}_{2}$, includes a patient population of size 1,859,536 and has relatively denser binary features from a different corpus.
Diseases of Interest. To investigate the utility of our proposed method, we consider using the generated synthetic EHR data to learn classifiers to predict six chronic diseases: type-II diabetes, chronic obstructive pulmonary disease (COPD), chronic kidney disease (CKD), asthma, hypertension heart and osteoarthritis. The prevalence of these diseases in each dataset is provided in Tabel 4 in Appendix A.
@@ 5.2 Baselines
Med-WGAN A number of GAN models (Choi et al., 2017b; Baowaly et al., 2018; Torfi and Fox, 2020a) have been proposed for realistic EHR generation. Torfi and Fox (2020a) utilizes convolutional neural networks, which is less applicable to most EHR generation as the correlated ICD codes are not in neighboring dimensions. Med-WGAN is selected as a baseline since it incorporates stable training techniques (Gulrajani et al., 2017; Hjelm et al., 2017) and has relatively robust performance.
EMR-WGAN Zhang et al. (2019) Different from other GAN models, which use an autoencoder to first transform the raw EHR data into a low-dimensional continuous vector, EMR-WGAN is directly trained on discrete EHR data.
EHRDiff Yuan et al. (2023) is the only diffusion model directly designed for synthesizing tabular EHR with an open-source codebase. As the code of other diffusion models (Yuan et al., 2023; Ceritli et al., 2023) for tabular EHRs are not available, we select EHRDiff as a baseline.
Refer to caption Refer to caption Refer to caption Refer to caption
@@ 5.3 Evaluation Metrics
Maximum Mean Discrepancy (MMD) is one of most common metrics to measure the difference between two distributions in distributional space. We compute the MMD for a set of synthetic data and a set of real test data. We employ a mixture of kernel-based methods (Li et al., 2017) to estimate MMD to improve its robustness. The detailed formula is given in Eq.(6) in Appendix A.4.
Downstream Prediction To evaluate the utility of the generated synthetic data, we evaluate the accuracy of classifiers trained to predict the diseases of interest mentioned above using synthetic data. We train the classifiers to predict the ICD code that corresponds to the disease of interest using all other available ICD codes as features. We train the classification model using synthetic data and evaluate its performance on real test data. We adopt the state-of-the-art robust classification model for tabular data given in (Ke et al., 2017). The most reliable classification model is one trained on real data; we use this as a benchmark to represent an upper bound for classification accuracy.
Membership Inference Risk (MIR) evaluates the risk that an attacker can infer the real samples used for training the generative model given generated synthetic EHR data or the model parameters. We consider the attack model MIR (Duan et al., 2023) proposed for diffusion models with continuous data. Following Yan et al. (2022), we evaluate MIR on discrete data. For each EHR in this set of real training and evaluation data, we calculate the minimum L2 distance with respect to the synthetic EHR data. The real EHR whose distance is smaller than a preset threshold is predicted as the training EHR. We report the prediction F1 score to demonstrate the performance of each model under membership inference risk.
@@ 5.4 Experiment Results
Table 1: Synthetic data utility. Disease prediction from ICD codes using the real dataset $\mathcal{D}_{1}$. AUPRC and AUROC are reported. AUPR and AUC in table are short for AUPRC and AUROC respectively. We use synthetic data of size 160K to train the classifier and 200K real test data samples to evaluate the different methods. 80% of the test data are bootstrapped 50 times to compute 95% confidence intervals (CI). The values of CI for all cases are between 0.001 and 0.003 and therefore not shown in table, provided on Appendix B.
Diabetes Asthma COPD CKD HTN-Heart Osteoarthritis
AUPR AUC AUPR AUC AUPR AUC AUPR AUC AUPR AUC AUPR AUC
Real Data 0.702  0.808 0.288  0.759 0.675  0.867 0.806  0.913 0.253  0.832 0.296  0.789
Med-WGAN 0.628  0.757 0.149  0.595 0.578  0.806 0.722  0.873 0.114  0.625 0.192  0.661
EMR-WGAN 0.656  0.770 0.193  0.642 0.603  0.815 0.753  0.885 0.151  0.686 0.219  0.689
EHRDiff 0.670  0.780 0.232  0.722 0.642  0.856 0.782  0.902 0.150  0.714 0.245  0.759
EHR-D3PM 0.693  0.801 0.263  0.748 0.655  0.860 0.796  0.908 0.229  0.821 0.278  0.782
@@
We first evaluate the learnt distributions using the MIMIC-III dataset in Fig. 1, dataset $\mathcal{D}_{1}$ in Fig. 3 and dataset $\mathcal{D}_{2}$ in Fig. 7 in Appendix B.1. We compare prevalence in synthetic data with prevalence in real data for each dimension. In Fig. 1, Fig. 3 and Fig. 7, we can see that the prevalence for our method EHR-D3PM aligns best with the real data. EHR-D3PM consistently has the highest Spearman correlation. We further observe that Med-WGAN, EMR-WGAN and EHRDiff fail to provide an unbiased estimation of the distribution in the low prevalence regime, which corresponds to rare conditions. This failure is mild when the dataset has dense features, as shown in Fig. 7 of $\mathcal{D}_{2}$, but is obvious when the dataset has sparse features, as shown in Fig. 3 of $\mathcal{D}_{1}$.
@@
Next we compare our method in terms of feature number per record, which is calculated by summing the ICD codes in each sample, using the MIMIC dataset-III in Fig. 2, dataset $\mathcal{D}_{1}$ in Fig. 4 and dataset $\mathcal{D}_{2}$ in Fig. 8 in Appendix B.1. We compare feature numbers for synthetic data with that of real data. As we see in Fig. 2, Fig. 4 and Fig. 8, Med-WGAN, EMR-WGAN and EHRDiff demonstrate poor performance in estimating the mode or the tail of the density. When the datasets are large, our method tends to provide a perfect estimation of the feature number for the real data.
@@
Finally we compare our method in terms of the CMD and MMD metrics. Tabel 2 shows that our EHR-D3PM significantly outperforms all baselines, particularly with the larger datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. This indicates our method can learn much better pairwise correlations between different feature dimensions. Table 2 also shows that the distribution of synthetic data generated by our method has the least discrepancy with real data distribution on all three datasets.
@@
Table 2: Additional fidelity metrics (CMD and MMD) and privacy metric (MIR) on MIMIC, dataset $\mathcal{D}_{1}$ and dataset $\mathcal{D}_{2}$. 95% confidence intervals are provided in Table 11 and Table 11 on Appendix B.
Fidelity CMD($\downarrow$) Fidelity MMD($\downarrow$) Privacy MIR($\downarrow$)
MIMIC         $\mathcal{D}_{1}$          $\mathcal{D}_{2}$ MIMIC       $\mathcal{D}_{1}$          $\mathcal{D}_{2}$ MIMIC       $\mathcal{D}_{1}$          $\mathcal{D}_{2}$
Med-WGAN 27.540    18.107    28.942 0.078      0.075      0.086 0.440      0.339      0.398
EMR-WGAN 26.658    11.869    21.438 0.053      0.018      0.024 0.456      0.358      0.415
EHRDiff 25.447    23.208    18.941 0.009      0.023      0.046 0.445      0.353      0.421
EHR-D3PM 21.128     7.692    10.255 0.003      0.012      0.019 0.432      0.344      0.406
@@
Utility We now apply our method to disease classification downstream tasks. Since MIMIC-III contains a much smaller patient population than the private datasets, which may not provide a valid test data benchmark for disease classification, we focus on the private datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}.$ In Table 4, we observe that the prevalence of most diseases is low in datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}.$ The training and test data set sizes are 160K and 200K. From Table 1, the average increase in AUPRC and AUROC over the strongest baseline (EHRDiff) is 3.90% and 2.57% respectively for dataset $\mathcal{D}_{1}$. From Table 5, the average increase in AUPRC and AUROC over the strongest baseline (EHRDiff) is 3.22% and 3.12% respectively for dataset $\mathcal{D}_{2}.$
@@
Privacy We evaluate the membership inference risk (MIR) of our method and other baselines on each dataset. Table 2 shows the MIRs of our method is lower than other baslines across all datasets, indicating our method has mild vulnerability to privacy risk compared to existing baselines. As there is a trade-off between privacy and fidelity, incorporating differential privacy to further reduce the privacy risk in diffusion models is an interesting direction, which is largely unexplored in discrete diffusion models.
@@
Figure 5: Synthetic data augmentation for disease classifications from ICD codes based on dataset $\mathcal{D}_{2}$. The size of the real source data for training the LGBM classifier is 5000, as indicated by the dashed purple line. We augment the source training data with synthetic data to train the LGBM classifier. "Uncond Samples" stands for the synthetic data generated by our unconditional sampler. Guided samples are synthetic data generated by our proposed guided sampler for each disease. To minimize noise from evaluation, we adopt 200K real test data to evaluate all experiments and report test AUROC for comparison. 80% of the test data are bootstrapped 50 times to compute 95% CI, which is added in shaded region.
@@ 5.5 Guided Generation
In this section, we apply our guided sampling method to generate conditional samples for different disease conditions. For each condition, we apply our guided sampler and generate a set of synthetic data. Table 3 lists the prevalence of four diseases in the real data $\mathcal{D}_{2}$. As we can see, the prevalence of diabetes, COPD, Asthma and Osteoarthritis are very low. The prevalence of samples generated by our unconditional sampling method matches the prevalence of each in the real datam, while synthetic samples generated by our guided sampler demonstrate a more balanced ratio.
Table 3: Prevalence of diseases in different sample groups. "Real" means real dataset $\mathcal{D}_{2}$. "Uncond" is short for samples generated by our unconditional method. "Guided" stands for samples generated by our guided sampler.
Diabetes COPD Asthma Osteoarthritis
Real 0.068 0.017 0.079 0.034
Uncond 0.069 0.015 0.073 0.036
Guided 0.202 0.042 0.161 0.137
In the following, we utilize synthetic samples to augment the training dataset when training downstream disease classifiers. The size of the real source data for training classifiers is 5000. We augment the original training data with data from three different groups: real data, synthetic data generated by our unconditional sampling method and our guided sampling method. We report AUROC in Fig. 5 and AUPRC in Fig. 9 in AppendixB.2 to evaluate accuracy. We can see that classifiers trained with synthetic data augmentation always improve the vanilla baseline (classifier trained with the original source data). We also observe that the data augmentation by guided sampling consistently outperform the data augmentation by unconditional sampling. It is interesting to observe that the data augmentation by guided sampling has consistently higher AUROC than real data augmentation. We observe this to be because the synthetic samples generated by guided sampling contain richer information in cases of diseases of low prevalence.

### v2

@@ 5.1 Experiment Setup
Datasets. Public Datasets MIMIC-III (Johnson et al., 2016) includes deidentified patient EHRs from hospital stays. For each patient’s EHR, we extract the diagnosis and procedure ICD-9 codes and truncate the codes to the first three digits. This dataset includes a patient population of size 46,520. Private Datasets We consider two private datasets of patient EHRs from a large healthcare institution. On $\mathcal{D}_{1}$, we follow the same process as MIMIC to extract ICD-10 codes. $\mathcal{D}_{1}$ includes a patient population of size 1,670,347 and has sparse binary features. The second dataset, denoted by $\mathcal{D}_{2}$, includes a patient population of size 1,859,536 and has relatively denser binary features from a different corpus. $\mathcal{D}_{2}$ contains diagnose codes (ICD), procedure codes (CPT) and medication codes (GEN). The total dimension is 2683.
Diseases of Interest. To investigate the utility of our proposed method, we consider using the generated synthetic EHR data to learn classifiers to predict six chronic diseases: type-II diabetes, chronic obstructive pulmonary disease (COPD), chronic kidney disease (CKD), asthma, hypertension heart and osteoarthritis. The prevalence of these diseases in each dataset is provided in Table 5 in Appendix A.
@@ 5.2 Baselines
Med-WGAN and EMR-WGAN A number of GAN models (Choi et al., 2017b; Baowaly et al., 2018; Torfi and Fox, 2020a) have been proposed for realistic EHR generation. Med-WGAN is selected as a baseline since it incorporates stable training techniques (Gulrajani et al., 2017; Hjelm et al., 2017) and has relatively robust performance. Different from other GAN models, which use an autoencoder to first transform the raw EHR data into a low-dimensional continuous vector, EMR-WGAN is directly trained on discrete EHR data.
TabDDPM and TabSyn We compare our model with two recent diffusion models, TabDDPM (Kotelnikov et al., 2023) and TabSyn (Zhang et al., 2023) for tabular data. When training TabSyn in first 100 dimensions of our datasets, TabSyn works well. But when training TabSyn in first 200 dimensions or all dimensions of our datasets, the performance of TabSyn degenerates significantly. One reason could be that the feature in EHRs are extremely sparse, which is challenging for learning in the first component of TabSyn. Therefore we leave the comparison with TabSyn on Appendix  B.1.
EHRDiff Yuan et al. (2023) is the only diffusion model directly designed for synthesizing tabular EHR with an open-source codebase. As the code of other diffusion models (Yuan et al., 2023; Ceritli et al., 2023; He et al., 2023a) for tabular EHRs are not available, we select EHRDiff as a baseline.
@@ 5.3 Evaluation Metrics
Maximum Mean Discrepancy (MMD) is one of most common metrics to measure the difference between two distributions in distributional space. We compute the MMD for a set of synthetic data and a set of real test data. We employ a mixture of kernel-based methods (Li et al., 2017) to estimate MMD to improve its robustness. The detailed formula is given in Eq. (6) on Appendix A.4.
Medical Concept Abundance Distance (MCAD) measures the synthetic EHR data distribution on the record level. MCAD metric quantifies the discrepancy between the empirical distribution (histogram) of the unique positive code number on synthetic data and the empirical distribution of the unique positive code number on real data.
Downstream Prediction To evaluate the utility of the generated data, we evaluate the accuracy of classifiers trained to predict the diseases of interest mentioned above using synthetic data. We train the classifiers to predict the ICD code that corresponds to the disease of interest using all other available ICD codes as features. We train the classification model using synthetic data and evaluate its performance on real test data. We adopt the state-of-the-art robust classification model for tabular data given in (Ke et al., 2017). The most reliable classification model is one trained on real data; we use this as a benchmark to represent an upper bound for classification accuracy.
Privacy Metrics Attribute inference risk (AIR) and membership inference risk (MIR) are employed to evaluate the vulnerability risks of our model. AIR measures the risk from the adversary ability to infer sensitive attributes of a targeted record when a subset of features are exposed to the attacker. AIR is calculated as the weighted sum of F1 scores of the inferences of other sensitive attributes. The number of exposed attributes for all experiments is defaulted as 256. MIR evaluates the risk that an attacker can infer the real samples used for training the generative model given generated EHR data or the model parameters. For each EHR in this set of real training and evaluation data, we calculate the minimum L2 distance with respect to the synthetic EHR data. The real EHR whose distance is smaller than a preset threshold is predicted as the training EHR. The predicted F1 score is computed to evaluate membership vulnerability risk. The threshold is set as 3 for all experiments. We use codebase of Yan et al. (2022) to compute both AIR and MIR scores.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Spearman corr = 0.93 (b) Spearman corr = 0.95 (c) Spearman corr = 0.89 (d) Spearman corr = 0.99
Refer to caption Refer to caption Refer to caption Refer to caption
(e) Spearman corr = 0.90 (f) Spearman corr = 0.94 (g) Spearman corr = 0.81 (h) Spearman corr = 0.98
Figure 1: Comparison of prevalence on synthetic data and real data $\mathcal{D}_{2}$ with ICD, CPT and GEN codes, where the total dimension is 2683. The second row represents the prevalence of the first row in the low data regime. The prevalence is computed on 200K samples. The dashed diagonal lines represent the perfect matching of code prevalence between synthetic data and real EHR data. Pearson correlations are very high for all methods and thus not used as a metric to compare different methods.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Density comparison of per-record feature number on synthetic data and real data $\mathcal{D}_{2}$ with ICD, CPT and GEN codes. The number of features per record is the sum of ICD codes present in each sample. The number of bins is 175, and the range of feature number values is (0, 175).
@@ 5.4 Experiment Results
We first evaluate the accuracy of marginal distributions on synthetic data in Figure 1 for dataset $\mathcal{D}_{2}$, Figure 5 for MIMIC and Figure 7 for $\mathcal{D}_{1}$ on Appendix B.1. We compare prevalence in synthetic data with prevalence in real data for each dimension. In Figure 1, Figure 5, and Figure 7, we can see that the prevalence for our method EHR-D3PM aligns best with the real data. EHR-D3PM consistently has the highest Spearman correlation. We further observe that EMR-WGAN and EHRDiff fail to provide an unbiased estimation of the distribution in the low prevalence regime, which corresponds to rare conditions. This failure is mild when the dataset has dense features, as shown in Figure 1 of $\mathcal{D}_{2}$, but is obvious when the dataset has sparse features, as shown in Figure 7 of $\mathcal{D}_{1}$.
@@
Next we visualize the histogram of feature number per record, which is calculated by summing the medical codes in each sample, on dataset $\mathcal{D}_{2}$ in Figure 2, MIMIC-III dataset in Figure 6, and dataset $\mathcal{D}_{1}$ in Figure 8 on Appendix B.1. We compare feature numbers for synthetic data with that of real data. As we see in Figure 6, Figure 8 and Figure 2, EMR-WGAN and EHRDiff demonstrate poor performance in estimating the mode or the tail of the density. When the datasets are large, our method tends to provide a perfect estimation of the feature number for the real data.
@@
Finally we compare our method with baselines in fidelity metrics, CMD, MMD and MCAD in Table 1. The results in CMD metric show that our EHR-D3PM significantly outperforms all baselines, particularly with the larger datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. This indicates our method can learn much better pairwise correlations between different feature dimensions. Table 1 also shows that the distribution of synthetic data learned by our method has the least discrepancy with real data distribution on all three datasets.
@@
Table 1: Fidelity metrics (CMD, MMD and MCAD) on MIMIC and real dataset $\mathcal{D}_{2}$. These metrics on real dataset $\mathcal{D}_{1}$ are provided on Appendix due to space limit.
Metrics CMD($\downarrow$) MMD($\downarrow$) MCAD($\downarrow$)
Datasets MIMIC                  $\mathcal{D}_{2}$ MIMIC                  $\mathcal{D}_{2}$ MIMIC                  $\mathcal{D}_{2}$
Med-WGAN 27.540 $\pm$ 0.628  28.942 $\pm$ 0.196 0.078 $\pm$ 0.0089  0.086 $\pm$ 0.013 0.1896 $\pm$ 0.0024  0.1944 $\pm$ 0.0017
EMR-WGAN 26.658 $\pm$ 0.639  21.438 $\pm$ 0.146 0.053 $\pm$ 0.0054  0.024 $\pm$ 0.004 0.1546 $\pm$ 0.0167  0.1572 $\pm$ 0.0015
TabDDPM 25.236 $\pm$ 0.684  16.728 $\pm$ 0.126 0.010 $\pm$ 0.0016  0.042 $\pm$ 0.008 0.1306 $\pm$ 0.0018  0.1587 $\pm$ 0.0018
EHRDiff 25.447 $\pm$ 0.485  18.941 $\pm$ 0.092 0.009 $\pm$ 0.0013  0.046 $\pm$ 0.005 0.1439 $\pm$ 0.0015  0.1764 $\pm$ 0.0016
EHR-D3PM 21.128 $\pm$ 0.393  10.255 $\pm$ 0.037 0.003 $\pm$ 0.0004  0.019 $\pm$ 0.002 0.1013 $\pm$ 0.0010  0.1081 $\pm$ 0.0009
T2D Asthma COPD CKD HTN-Heart Osteoarthritis
Med-WGAN 0.924 $\pm$ 0.001 0.819 $\pm$ 0.002 0.853 $\pm$ 0.003 0.835 $\pm$ 0.004 0.500 $\pm$ 0.001 0.820 $\pm$ 0.003
TabPPDM 0.945$\pm$ 0.003 0.846$\pm$ 0.003 0.931$\pm$ 0.004 0.915$\pm$ 0.009 0.837$\pm$ 0.008 0.868 $\pm$ 0.005
@@
Utility We now apply our method to disease classification downstream tasks. Since MIMIC-III contains a much smaller patient population than the private datasets, which may not provide a valid test data benchmark for disease classification, we focus on the private datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}.$ In Table 5, we observe that the prevalence of most diseases is low in datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}.$ The training and test data set sizes are 160K and 200K. From Table 2 and Table 8 on Appendix  B.1, the absolute average increase in AUPRC and AUROC over baselines (EHRDiff and TabDDPM) is 3.22% and 3.12% respectively for dataset $\mathcal{D}_{2}.$ From Table 6 and Table 7 on Appendix  B.1, the absolute average increase in AUPRC and AUROC over baselines (EHRDiff and TabDDPM) is 3.90% and 2.57% respectively for dataset $\mathcal{D}_{1}$. The improvement in downstream prediction for our model benefits from the best correlation between different features our model has learned.
@@
Privacy In Table 3, we evaluate the attribute inference risk (AIR) and membership inference risk (MIR) of our model and other baselines. It shows that our model has relatively low risks on both AIR and MIR. This indicates our method has mild vulnerability to privacy risk compared to existing baselines. Generally, there is a trade-off between privacy and fidelity for generative models. In particular, when a generative model fails to learn any information of the data distribution, the AIR and MIR risk scores are expected to be the lowest. Therefore, incorporating differential privacy to further reduce the privacy risk in diffusion models is an interesting direction, which is largely unexplored in discrete diffusion models.
Table 3: Privacy metrics (AIR and MIR) on MIMIC, dataset $\mathcal{D}_{1}$ and dataset $\mathcal{D}_{2}$.
Attribute Inference Risk($\downarrow$) Membership Inference Risk($\downarrow$)
MIMIC                $\mathcal{D}_{1}$                   $\mathcal{D}_{2}$ MIMIC                $\mathcal{D}_{1}$                   $\mathcal{D}_{2}$
Med-WGAN 0.019$\pm$0.0015  0.071$\pm$0.0056  0.080$\pm$0.0056 0.440$\pm$0.0034  0.339$\pm$0.0018  0.398$\pm$0.0019
EMR-WGAN 0.058$\pm$0.0027  0.116$\pm$0.0084  0.132$\pm$0.0089 0.456$\pm$0.0035  0.358$\pm$0.0017  0.415$\pm$0.0020
TabDDPM 0.021$\pm$0.0013  0.085$\pm$0.0091  0.093$\pm$0.0062 0.462$\pm$0.0037 0.351$\pm$0.0018  0.415$\pm$0.0020
EHRDiff 0.022$\pm$0.0016  0.077$\pm$0.0042  0.089$\pm$0.0054 0.445$\pm$0.0034  0.353$\pm$0.0015  0.421$\pm$0.0019
EHR-D3PM 0.020$\pm$0.0014  0.068$\pm$0.0046  0.078$\pm$0.0048 0.432$\pm$0.0034  0.344$\pm$0.0016  0.406$\pm$0.0018
@@
Figure 3: Synthetic data augmentation for disease classifications from ICD codes based on dataset $\mathcal{D}_{2}$. The size of the real source data for training the LGBM classifier is 5000, as indicated by the dashed purple line. We augment the source training data with synthetic data to train the LGBM classifier. "Uncond Samples" stands for the synthetic data generated by our unconditional sampler. Guided samples are synthetic data generated by our proposed guided sampler for each disease. To minimize noise from evaluation, we adopt 200K real test data to evaluate all experiments and report test AUROC for comparison. 80% of the test data are bootstrapped 50 times to compute 95% CI, which is visualized by the shaded region around each line.
@@ 5.5 Guided Generation
In this section, we apply our guided sampling method to generate conditional samples for different disease conditions. For each condition, we apply our guided sampler and generate a set of synthetic data. Table 4 lists the prevalence of four diseases in the real data $\mathcal{D}_{2}$. As we can see, the prevalence of diabetes, COPD, Asthma and Osteoarthritis are very low. The prevalence of samples generated by our unconditional sampling method matches the prevalence of each in the real data, while synthetic samples generated by our guided sampler demonstrate a more balanced ratio.
Table 4: Prevalence of diseases in different sample groups. "Real" means real dataset $\mathcal{D}_{2}$. "Unconditional" denotes samples generated by our unconditional method. "Guided" stands for samples generated by our guided sampler.
Diabetes Asthma COPD Osteoarthritis
Real 0.068 $\pm$ 0.002 0.079 $\pm$ 0.002 0.017 $\pm$ 0.000 0.034 $\pm$ 0.002
Unconditional 0.069 $\pm$ 0.003 0.073 $\pm$ 0.001 0.015 $\pm$ 0.001 0.036 $\pm$ 0.001
Guided 0.202 $\pm$ 0.004 0.161 $\pm$ 0.003 0.042 $\pm$ 0.001 0.137 $\pm$ 0.002
In the following, we utilize synthetic samples to augment the training dataset when training downstream disease classifiers. The size of the real source data for training classifiers is 5000. We augment the original training data with data from three different groups: real data, synthetic data generated by our unconditional sampling method and our guided sampling method. We report AUROC in Figure 3 and AUPRC in Figure 9 on Appendix B.2 to evaluate accuracy. We can see that classifiers trained with synthetic data augmentation always improve the vanilla baseline (classifier trained with the original source data). We also observe that the data augmentation by guided sampling consistently outperform the data augmentation by unconditional sampling. It is interesting to observe that the data augmentation by guided sampling has consistently higher AUROC than real data augmentation. We observe this to be because the synthetic samples generated by guided sampling contain richer information in cases of diseases of low prevalence.